Please help with these equations: calculate the products; (-i)^4×3i^3
Answer by lwsshak3(11628)

calculate the products; (-i)^4×3i^3
i=√-1
i^2=-1
i^3=-i
i^4=1
(-i)^4×3i^3=1*-3i=-3i